OpenStudy (anonymous):

Suppose n = pq where p and q is a twin prime pair (i.e. they di er by 2). Explain how this information can help us factor n quickly. Use your method to factor 52012943.

OpenStudy (anonymous):

twin prime.. q = p+2

OpenStudy (anonymous):

n = (p)(p+2)=p^2+2p p^2+2p-n=0 find p by quadratic

OpenStudy (anonymous):

I tried but the values i got for p were not prime

OpenStudy (anonymous):

\[ p=\frac{-2\pm\sqrt{4+4(52012943)}}{2}\\ p=\frac{-2\pm2\sqrt{1+(52012943)}}{2}\\ p=-1\pm\sqrt{52012944}\quad=-1+7212,\quad-1-7212 \] reject the -ve

OpenStudy (anonymous):

thanks btw how do you put the root sign